OpenStudy (lgbasallote):

find a solution for \[(y - \cos^2 x)dx + \cos x dy = 0\]

OpenStudy (lgbasallote):

do i divide by cosxdx?

OpenStudy (anonymous):

i'm trying to figure out which method would work best lol

OpenStudy (anonymous):

you might be able to ge this into a linear

OpenStudy (anonymous):

oh wait maybe y=ux

OpenStudy (anonymous):

y=ux dy=udx+xdu

OpenStudy (anonymous):

\[\frac{dy}{dx}+\frac{y}{cosx}=cosx\]

OpenStudy (lgbasallote):

homogeneous??? o.O

OpenStudy (lgbasallote):

i hate that thing lol @THE_PROPHET so you divided by cosxdx right?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

ok thanks